OpenStudy (anonymous):

Can someone please help me, I have been waiting for help for almost an hour. Please help

OpenStudy (anonymous):

What's your question?

OpenStudy (anonymous):

In 2005, some data was taken on the number of women in certain age groups who gave birth.? Age of Women 22 27 32 Number of women per 102 116 96 1000 who gave birth a)determain f(x)=a(x-h)^2+k,so that f model this data(not ax^2+bx+c)(Use trial and error, and find values of a, h, and k to the tenths place to guarantee that your graph comes close enough to “go through” all 3 points on the table when on “ZOOM STAT” [as boxes in the TI-83/84].) (b) Estimate the number of women per 1000 who give birth at age 35 according to your model (c) Where does it look like f (x)=0 ? (You will likely need to expand your window, especially since you used “ZoomStat”.) What does this mean? Do you believe this? Is f(x)a perfect model of what really happens? (answer all 3 questions)

OpenStudy (anonymous):

Shoot, I don't have a graphing calculator with me. Sorry! I could help you do it by hand, but it'll take a lot longer.

OpenStudy (anonymous):

thats ok, I have an exam on this so i would like to learn it by hand or calculator , if you can?

OpenStudy (anonymous):

Alright. Your coordinates are (22, 102), (27, 116), and (32, 96). Your general form is going to be y = ax^2 + bx + c.

OpenStudy (anonymous):

he said Not ax^2 +bx+c

OpenStudy (anonymous):

Yeah, but this is the only way I know how to do it by hand. Your graphing calculator is set up to put it into the other form.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

The first step is to determine your equations. Take the first point (22, 102). 22 is the x-coordinate, so you're going to substitute 22 for x. The y-coordinate is 102. Substitute that for y. You get \[102 = a*22^2+b*22+c\]

OpenStudy (anonymous):

Okay with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

*What level of math are you at?

OpenStudy (anonymous):

95

OpenStudy (anonymous):

Roughly Intermediate Algebra?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So this equation simplifies this way 102 = 484a + 22b + c

OpenStudy (anonymous):

Try finding the equation for the second point.

OpenStudy (anonymous):

116=729a+27b+c ?

OpenStudy (anonymous):

Yep, good :). Third one?

OpenStudy (anonymous):

96=1024a+32b+c

OpenStudy (anonymous):

Good. Now we have our three equations: 102 = 484a + 22b + c 116 = 729a + 27b + c 96 = 1024a + 32b + c Let's take the first equation. 102 = 484a +22b + c. Solve that for c. What do you get?

OpenStudy (anonymous):

can you show me the first one

OpenStudy (anonymous):

Sure. We want c, so we have to get rid of 484a and 22b. Subtract 484a from both sides to get -484a + 102 = 22b + c Okay with that step?

OpenStudy (anonymous):

would you also subtract 48 from the other side

OpenStudy (anonymous):

484a? Yes, but I already took it off.

OpenStudy (anonymous):

102 - 484a = 484a - 484a +22b + c

OpenStudy (anonymous):

-484a + 102 = 22b + c

OpenStudy (anonymous):

ok i’m sorry i am soo lost

OpenStudy (anonymous):

when i did it using the Ti-83 i got a=-.68 h-26.56 k=116.13

OpenStudy (anonymous):

No worries. We start with this: 102 = 484a + 22b + c Subtract 484a from both sides: 102 = 484a + 22b + c -484a -484a -484a +102 = 0a + 22b + c -484a +102 = 22b +c

OpenStudy (anonymous):

OH, missed your last message.

OpenStudy (anonymous):

That makes it much easier.

OpenStudy (anonymous):

So you have f(x) = a(x-h)^2 + k Substitute your values: f(x) = -0.68(x - 26.56)^2 + 116.13

OpenStudy (anonymous):

That function is a model for your data.

OpenStudy (anonymous):

where i am stuck at is part b: Estimate the number of women per 1000 who give birth at age 35 according to your model.

OpenStudy (anonymous):

What does x represent in your model?

OpenStudy (anonymous):

If you understand that, then the problem will be easy.

OpenStudy (anonymous):

age of women i think

OpenStudy (anonymous):

Yep. Which is 35 in this case. Just plug that into x and whatever comes out will be your answer.

OpenStudy (anonymous):

f(35) = -0.68(35 - 26.56)^2 + 116.13

OpenStudy (anonymous):

67.691152?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

ok so what other part of the question is what does this mean? do you belive this? Is f(x) a perfect model of what really happens?

OpenStudy (anonymous):

Well, the model they chose for this is a quadratic equation, which doesn't really make sense. There's not really any reason to suspect that it would be a quadratic equation, but it would probably approximate it pretty well for a small area of the data.

OpenStudy (anonymous):

Also, it will dip below zero eventually, which doesn't make sense either.

OpenStudy (anonymous):

where does it look like f(x)=0

OpenStudy (anonymous):

That's where it hits the x-axis. Zoom out until you see it hit the axis.

OpenStudy (anonymous):

thank you so much, you made it clear to understand the steps

OpenStudy (anonymous):

No problem :) . Have a good night.

OpenStudy (anonymous):

thanks you too